Statement of Cost of Goods Manufactured from Percent Relationships
Information about NuWay Products Company for the year ending December 31, 2010, follows:

  • Sales equal $500,000.
  • Direct materials used total $51,000.
  • Manufacturing overhead is 150 percent of direct labor dollars.
  • The beginning inventory of finished goods is 20 percent of the cost of goods sold.
  • The ending inventory of finished goods is twice the beginning inventory.
  • The gross profit is 20 percent of sales.
  • There is no beginning or ending work-in-process.

Prepare a statement of cost of goods manufactured for 2010. (Hint: Prepare an analysis of changes in Finished Goods Inventory.)

Do not use negative signs with any of your answers below.

NuWay Products Company
Statement of Cost of Goods Manufactured
For the Year Ending December 31, 2010
Current manufacturing costs:                    
Direct materials
Direct labor
Manufacturing overhead
Beginning work-in-process
Total costs in process
Ending work-in-process
Cost of goods manufactured
Analysis of Finished Goods Inventory:
Finished goods, 1/1/10
Cost of goods manufactured
Total goods available for sale
Finished goods, 12/31/10
Cost of goods sold                     
Answer #1

Answer :

Preparation of statement of cost of goods manufactured for 2010 :

NuWay Products Company
Statement of Cost of Goods Manufactured
For the Year Ending December 31, 2010
Current Manufacturing Costs :

Direct Materials

$51,000

Direct Labor

171,600

Manufacturing Overhead

257,400 $480,000
Plus : Beginning Work in Process 0
Total Costs in process $480,000
Less : Ending Work in Process 0
Cost of goods manufactured $480,000
Analysis of finished goods inventory :
Finished goods 1/1/10 $80,000
Plus : Cost of goods manufactured 480,000
Total goods available for sale 560,000
Less : Finished goods, 12/31/10 (160,000)
Cost of goods sold $400,000

Working Notes :-

1. Cost of goods sold = Sales - Gross profit

= $ 500,000 - 20 % = $ 400,000

2. Beginning inventory of finished goods = 20 % of Cost of goods sold

= $400,000 × 20 % = $ 80,000

3. Ending inventory of finished goods = 2 × Beginning inventory of finished goods

= $ 2 × 80,000 = $ 160,000

4. Total goods available for sale = Cost of goods sold + Ending inventory of finished goods

= $ 400,000 + 160,000 = $ 560,000

5. Cost of goods manufactured = Total goods available for sale - Beginning inventory of finished goods

= $ 560,000 - 80,000 = $ 480,000

6. Calculation of Direct Labor and Manufacturing overhead:

Let direct labor be 'x'. Then, Manufacturing overhead will be 150 % of direct labor i.e, ' 1.5 x '

Total Manufacturing costs incurred during the year = Direct materials + Direct labor + Manufacturing overhead

$ 480,000 = $ 51,000 + x + 1.5 x

$ 480,000 - $ 51,000 = 2.5 x

$ 429,000 ÷ 2.5 = x

X = $ 171,600 = Direct Labor and ,

Manufacturing Overhead = 1.5 × 171,600 =$257,400